I have a cylinder 48" tall & 30" in diameter. How many gallons will it hold?: I have a cylinder 48" tall & 30" in diameter. How many gallons will it hold?
Answer by jim_thompson5910(9927)

Remember, if the diameter is 30, then the radius is 15

The volume of a cylinder can be found by the formula
V = pi*R^2*H

Now plug in R=15 and H=48

V = pi*15^2*48 = 33929.20062


So it can hold about 33,929.2 gallons